Two years ago a father was four times as old as his son. Three years from now the father will be only three times as old as the son. How old is each now?

Answer by ankor@dixie-net.com(22740)

Let f = father's present age
Let s = son's present age
:
Just write an equation for what it says;
:
"Two years ago a father was four times as old as his son."
f - 2 = 4(s - 2)
f - 2 = 4s - 8
f = 4s - 8 + 2
f = 4s - 6
:
" Three years from now the father will be only three times as old as the son."
f + 3 = 3(s + 3)
f + 3 = 3s + 9
f = 3s + 9 - 3
f = 3s + 6
:
Replace f with (4s-6) from the 1st statement
4s - 6 = 3s + 6
4s - 3s = 6 + 6
s = 12 yrs is the son's age
:
I'll let you find the father's age, using one of the equations, then check your
solutions in the other equation